[figure: 9.7 cm]

OVERCRITICAL MOTION OF BLOCH DOMAIN WALL IN

ULTRATHIN FERROMAGNETIC FILMS

V.G. Baryakhtar^{a} V.V. Tarasenko^{b} and S.V. Tarasenko^{b} aInstitute of Magnetism, 252142 Kiev, Vernadsky Prospekt 36b, Ukraine

bInstitute of Physics, Bialystok University, 15-424 Bialystok, 41 Lipowa, Poland

The overcritical domain wall (DW) dynamics in ultrathin ferromagnetic film without magnetic losses is considered. The overcritical motion occurs when DW energy exceeds critical Walker value and oscillatory processes become important. In overcritical region the DW energy is defined by two parameters: a) population number (n) of the lowest oscillatory level and b) angle (phi) of the magnetic moment declination out of the DW plane. The DW velocity is defined by this population number and should be calculated with the help of the Goldstone and Adler theorems. The Lagrangian maximization procedure with respect to phi at a given energy allows to find phi, n and DW velocity as a function of energy. In ideal film without magnetic losses the DW velocity in the overcritical region does not depend on time. Inhomogeneties of the film lead to the velocity time oscillations as well as magnetic losses. The stability region for the overcritical motion is also defined.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : V.V. Tarasenko

Address for correspondence : Institute of Physics, 15-424 Bialystok, 41 Lipowa, Poland

Email address : taraserg@cksr.ac.bialystok.pl